OpenStudy (anonymous):

If f(x) = 1/(x-2), (f circle g)'(1) = 6 and g'(1) = -1, then g(1) =

OpenStudy (anonymous):

Here I'll post what I've tried...

OpenStudy (anonymous):

I did \[f'(g(1)) * g'(1) = 6\] \[-6 = f'(g(1))\] \[f'(x) = -1/(x-2)^2\] \[-6 = -1/(x-2)^2\] \[(x-2)^2 = 1/6\] \[x = 2 +-\sqrt{6}/6\] and g(1) = x

ganeshie8 (ganeshie8):

Looks good!

OpenStudy (anonymous):

It does? But it doesn't match any of my answer choices

OpenStudy (anonymous):

They are: a. -7 b. -5 c. 5 d. 7 e. 8

jimthompson5910 (jim_thompson5910):

I did it a slightly different way, but I'm getting the same result f(x) = 1/(x-2) f(g(x)) = 1/(g(x)-2) (f o g)(x) = 1/(g(x)-2) (f o g)'(x) = -(g'(x))/(g(x)-2)^2 (f o g)'(1) = -(g'(1))/(g(1)-2)^2 6 = -(-1)/(z-2)^2 ... let z = g(1) 6 = 1/(z-2)^2 6(z-2)^2 = 1 (z-2)^2 = 1/6 z-2 = +-sqrt(1/6) z-2 = +-sqrt(6)/6 z = 2+-sqrt(6)/6 g(1) = 2+-sqrt(6)/6

OpenStudy (anonymous):

Then I think they put the wrong answer choices. It's from a practice AP exam so there could be a mistake in it

ganeshie8 (ganeshie8):

unique answer is not possible because of the quadratic expression in the denominator of f'(x) so most likely there is some mistake

OpenStudy (anonymous):

That's true. I'll just have to ask my teacher when I come back from break

OpenStudy (anonymous):

Thank you for helping guys!